# Disordered-DABS: A Benchmark for Dynamic Aspect-Based Summarization in Disordered Texts

### v1

@@ 1 Introduction
Our paper addresses these challenges by presenting a novel and efficient method to modify existing summarization datasets, making them suitable for dynamic aspect-based summarization in disordered texts. We use this method to introduce the Disordered-DABS dataset [^2] The data is available on Dataverse, specifically curated for this demanding summarization task. Comprising 255,286 aspect-oriented summaries, this dataset is segmented into training, validation, and testing partitions. Its robustness is validated through comprehensive human evaluations. Our examination of various baseline models highlights the complexity of this task, revealing that even advanced large language models like GPT-3.5 face significant challenges in dealing with the nuances of Disordered-DABS. These observations underscore the intricacy inherent in this form of summarization, highlighting the necessity for more specialized datasets tailored to its unique requirements.
@@ 2 Related Work
Further advancing this evolution is DABS, which diverges from traditional ABS by not limiting aspects to predefined categories. Innovations in DABS, such as AnyAspect Tan et al. (2020) and ENTSUMv2 Mehra et al. (2023), have focused on utilizing entities as aspects, with AnyAspect summarizing sentences around named entities and ENTSUMv2 employing human-annotated, entity-specific summaries. Other datasets like OASUM Yang et al. (2023b) and OPENASP Amar et al. (2023a) explore conceptual labels or sub-topics as aspects, with OASUM leveraging Wikipedia subtitles and OPENASP adopting a multi-document DABS settings. A comparative overview of these datasets is presented in Table 1.
@@
OpenASP(Amar et al., 2023a) News manual 1,310
ENTSUMV2(Maddela et al., 2022) News automaic 2,788
@@ 3 Dynamic Aspect-Based Summarization
Despite substantial efforts in QFS and ABS, the definition of “aspect” remains ambiguous. It ranges from different topics Frermann and Klementiev (2019); Maddela et al. (2022); Tan et al. (2020) to specific elements within a topic Hayashi et al. (2021); Ahuja et al. (2022); Meng et al. (2021). Some studies use concise phrases for aspects Amar et al. (2023a); Angelidis et al. (2021), while others allow longer sentences for complex information needs Zhong et al. (2022, 2021). Given this variability in aspect definition, our Disordered-DABS incorporates two distinct sub-datasets: one emphasizing aspect diversity across varied topics, and another focusing on different aspects of a single event, providing a comprehensive benchmark for these divergent aspect interpretations.
Building on previous DABS research Tan et al. (2020); Yang et al. (2023b), our task is defined as generating multiple aspect-based summaries from a document $D$ containing $n$ topics or aspects. Unlike traditional baselines that pair a document with a specific aspect for aspect-based summary generation Amar et al. (2023a), our task requires the model to autonomously generate multiple summaries without predefined aspects. Additionally, to challenge the model’s ability to handle non-coherent texts, we shuffle the sentences within $D$, following Frermann and Klementiev (2019), disrupting the natural flow and requiring the model to cluster information from the entire document.
@@ 3.1 Automatic Evaluation Metrics
Automatic evaluation metrics are vital for the efficient development and evaluation of models for the ABS and DABS tasks. Our evaluation focuses on two primary factors: the capability to identify aspect-based information and the overall quality of the generated summaries. We measure the aspect identification accuracy using the absolute difference in aspects between the reference and generated summaries (#$\#AbsAspDiff{}$). For assessing summary quality, we employ the standard summary quality metrics Rouge-1/2/L Lin (2004) and BERTScore Zhang et al. (2019).
To compare reference and generated summaries effectively, we pad the summaries to equalize the number of aspects. This approach penalizes models that either miss aspects or generate excessive ones. The optimal pairing of reference and generated summaries ($<\hat{S},S>$) is determined by maximizing the mean performance across all metrics, as defined by:
$<\hat{S},S>=\text{argmax}\left(\sum_{\hat{s}\in\hat{S},s\in S}\left(\sum_{m\in M}m(\widehat{s},{s})\right)\right)$ (1)
@@ 5 Baseline Models
Both the Cluster and Keyword methods employ BERTopic Egger and Yu (2022) for sentence clustering and keyword generation, taking advantage of its capabilities in unsupervised clustering and topic modeling. The hyperparameters of BERTopic models are fine-tuned to align the number of references and generated summaries in validation datasets, with further details provided in Appendix B.4. Concurrently, the Prompting method utilizes prompts to direct large-language models, such as GPT-3.5, in automatically aggregating aspect-based information and crafting summaries.
@@ 5.1 Clustering-Based-Summarization
The Cluster method, inspired by previous work Hayashi et al. (2021); Amar et al. (2023b), employs BERTopic to cluster sentences based on their respective aspects. Unlike their supervised approach, we utilize an unsupervised model within BERTopic to handle dynamic aspects with varying numbers and contents. Post-clustering, an abstractive summarization model based on Longformer-Encode-DecoderBeltagy et al. (2020) is used for generating summaries. This summarization model is fine-tuned using original training sets formatted as one-document-to-one-summary, which may potentially inflate its performance.
@@ 6.2 Human Evaluations
D-WikiHow Prompting 4.55 (0.57) 3.58 (1.09) 4.39 (0.72) 4.10 (1.14) 4.58 (0.81) 1.00 (0.00)
Keyword 2.16 (1.29) 2.58 (0.72) 2.48 (1.03) 1.90 (0.94) 1.97 (1.35) 2.68 (0.48)
@@
Notably, Prompting was consistently rated highest for “Fluency” and 11Aspect Quality.” This is in contrast to automatic evaluations, where Prompting was penalized for generating fewer aspects. Further inquiry with judges revealed that Prompting tends to fragment one aspect into several, accounting for this discrepancy. Additionally, annotators’ preference for a higher number of aspects in D-WikiHow, as opposed to reference standards, highlights the subjective nature of determining the optimal number of aspects, especially in nuanced aspect differentiation.
@@ 7.2 The Impact of Disorder Texts
Table 7: The performance of baseline models on the D-CnnDM and D-WikiHow sub-datasets, noting the effects of organized settings in parentheses. For the ’#AbsAspDiff’ metric, a negative value signifies an improvement, whereas the opposite is true for the remaining four metrics.

### v2

@@ 1 Introduction
Our paper addresses these challenges by presenting a novel and efficient method to modify existing summarization datasets, making them suitable for dynamic aspect-based summarization in disordered texts. We use this method to introduce the Disordered-DABS dataset [^2] The data is available on Github, specifically curated for this demanding summarization task. Comprising 255,286 aspect-oriented summaries, this dataset is segmented into training, validation, and testing partitions. Its robustness is validated through comprehensive human evaluations. Our examination of various baseline models highlights the complexity of this task, revealing that even advanced large language models like GPT-3.5 face significant challenges in dealing with the nuances of Disordered-DABS. These observations underscore the intricacy inherent in this form of summarization, highlighting the necessity for more specialized datasets tailored to its unique requirements.
@@ 2 Related Work
Further advancing this evolution is DABS, which diverges from traditional ABS by not limiting aspects to predefined categories. Innovations in DABS, such as AnyAspect Tan et al. (2020) and ENTSUMv2 Mehra et al. (2023), have focused on utilizing entities as aspects, with AnyAspect summarizing sentences around named entities and ENTSUMv2 employing human-annotated, entity-specific summaries. Other datasets like OASUM Yang et al. (2023b) and OPENASP Amar et al. (2023) explore conceptual labels or sub-topics as aspects, with OASUM leveraging Wikipedia subtitles and OPENASP adopting a multi-document DABS settings. A comparative overview of these datasets is presented in Table 1.
@@
OpenASP(Amar et al., 2023) News manual 1,310
ENTSUMV2(Maddela et al., 2022) News automatic 2,788
@@ 3 Dynamic Aspect-Based Summarization
Despite substantial efforts in QFS and ABS, the definition of “aspect” remains ambiguous. It ranges from different topics Frermann and Klementiev (2019); Maddela et al. (2022); Tan et al. (2020) to specific elements within a topic Hayashi et al. (2021); Ahuja et al. (2022); Meng et al. (2021). Some studies use concise phrases for aspects Amar et al. (2023); Angelidis et al. (2021), while others allow longer sentences for complex information needs Zhong et al. (2022, 2021). Given this variability in aspect definition, our Disordered-DABS incorporates two distinct sub-datasets: one emphasizing aspect diversity across varied topics, and another focusing on different aspects of a single event, providing a comprehensive benchmark for these divergent aspect interpretations.
Building on previous DABS research Tan et al. (2020); Yang et al. (2023b), our task is defined as generating multiple aspect-based summaries from a document $D$ containing $n$ topics or aspects. Unlike traditional baselines that pair a document with a specific aspect for aspect-based summary generation Amar et al. (2023), our task requires the model to autonomously generate multiple summaries without predefined aspects. Additionally, to challenge the model’s ability to handle non-coherent texts, we shuffle the sentences within $D$, following Frermann and Klementiev (2019), disrupting the natural flow and requiring the model to cluster information from the entire document.
@@ 3.1 Automatic Evaluation Metrics
Automatic evaluation metrics are vital for the efficient development and evaluation of models for the ABS and DABS tasks. Our evaluation focuses on two primary factors: the capability to identify aspect-based information and the overall quality of the generated summaries. We measure the aspect identification accuracy using the absolute difference in aspects between the reference and generated summaries (#AbsAspDiff). For assessing summary quality, we employ the standard summary quality metrics Rouge-1/2/L Lin (2004) and BERTScore Zhang et al. (2019).
To compare reference and generated summaries effectively, we pad the summaries to equalize the number of aspects. This approach penalizes models that either miss aspects or generate excessive ones. The optimal pairing of reference and generated summaries ($\langle\hat{S},S\rangle$) is determined by maximizing the mean performance across all metrics, as defined by:
$\langle\hat{S},S\rangle=\text{argmax}\left(\sum_{\hat{s}\in\hat{S},s\in S}\left(\sum_{m\in M}m(\widehat{s},{s})\right)\right)$ (1)
@@ 5 Baseline Models
Both the Cluster and Keyword methods employ BERTopic Egger and Yu (2022) for sentence clustering and keyword generation, taking advantage of its capabilities in unsupervised clustering and topic modeling. The hyperparameters of BERTopic models are fine-tuned to align the number of references and generated summaries in validation datasets, with further details provided in Appendix B.4. Concurrently, the Prompting method utilizes prompts to direct large-language models, such as GPT-3.5, in automatically aggregating aspect-based information and crafting summaries. Based on previous study Guo and Vosoughi (2023), we control the length of the generated summaries of all models for mitigating the influence of the generated summary length.
@@ 5.1 Clustering-Based-Summarization
The Cluster method, inspired by previous work Hayashi et al. (2021); Amar et al. (2023), employs BERTopic to cluster sentences based on their respective aspects. Unlike their supervised approach, we utilize an unsupervised model within BERTopic to handle dynamic aspects with varying numbers and contents. Post-clustering, an abstractive summarization model based on Longformer-Encode-Decoder Beltagy et al. (2020) is used for generating summaries. This summarization model is fine-tuned using original training sets formatted as one-document-to-one-summary, which may potentially inflate its performance.
@@ 6.2 Human Evaluations
Prompting 3.21 (1.28) 2.71 (0.75) 4.29 (1.12) 3.12 (0.80) 4.25 (1.07) 2.08 (0.88)
D-WikiHow Keyword 2.16 (1.29) 2.58 (0.72) 2.48 (1.03) 1.90 (0.94) 1.97 (1.35) 2.68 (0.48)
@@
Notably, Prompting was consistently rated highest for “Fluency” and Aspect Quality.” This is in contrast to automatic evaluations, where Prompting was penalized for generating fewer aspects. Further inquiry with judges revealed that Prompting tends to fragment one aspect into several, accounting for this discrepancy. Additionally, annotators’ preference for a higher number of aspects in D-WikiHow, as opposed to reference standards, highlights the subjective nature of determining the optimal number of aspects, especially in nuanced aspect differentiation.
@@ 7.2 The Impact of Disorder Texts
Table 7: The performance of baseline models on the D-CnnDM and D-WikiHow sub-datasets, noting the effects of organized settings in parentheses. For the “#AbsAspDiff” metric, a negative value signifies an improvement, whereas the opposite is true for the remaining four metrics.